How much power is required to make a 1400 kg car accelerate from 0 ms-1 to 20+your birth date ms-1 in 4.5 seconds?


1
Expert's answer
2021-03-23T11:11:33-0400

Assume that your birth date 15.05.2005


"W=KE=\frac{mv^2}{2}=\frac{1400\cdot (20+15)^2}{2}=857500\ (J)"


"P=W/t=857500/4.5\approx191\ (kW)" . Answer